### v1

# OptiRepair: Closed-Loop Diagnosis and Repair of Supply Chain Optimization Models with LLM Agents

Abstract
Supply chain optimization models frequently become infeasible because of modeling errors. Diagnosis and repair require scarce OR expertise: analysts must interpret solver diagnostics, trace root causes across echelons, and fix formulations without sacrificing operational soundness. Whether AI agents can perform this task remains untested. OptiRepair decomposes this task into two phases: a domain-agnostic feasibility phase that iteratively repairs any LP using IIS-guided diagnosis, and a domain-specific validation phase that enforces five rationality checks grounded in inventory theory. We test 22 API models from seven families on 976 multi-echelon supply chain problems and train two 8B-parameter models with self-taught reasoning and solver-verified rewards. The trained models reach 81.7% Rational Recovery Rate (RRR)—the fraction of problems resolved to both feasibility and operational rationality—versus 42.2% for the best API model and 21.3% on average. The gap concentrates in Phase 1 repair, where API models average 27.6% recovery rate versus 97.2% for trained models. Two gaps separate current AI from reliable model repair: solver interaction, as API models restore only 27.6% of infeasible formulations; and operational rationale, as roughly one in four feasible repairs violate supply chain theory. Each gap requires a different intervention—targeted training closes the solver interaction gap, while explicit specification as solver-verifiable checks closes the rationality gap. For organizations adopting AI in operational planning, formalizing what “rational” means in their context is the higher-return investment.
Keywords: supply chain optimization; large language models; model diagnosis and repair; operational rationality; reinforcement learning
1 Introduction
Consider a supply chain analyst building a multi-echelon inventory optimization model in Gurobi. The model spans five echelons and 24 planning periods, with over 500 constraints encoding demand propagation, inventory balance, lead-time coupling, and production capacity. After submitting the formulation, the solver returns Infeasible. The analyst must determine which constraints conflict, understand why (a misspecified lead time? an incorrect capacity limit?), and repair the formulation so that the solution is not merely feasible but operationally sound. This iterative diagnosis-and-repair loop—IIS analysis, interpretation, model modification, re-solving, and operational validation—consumes hours of expert time and recurs with every model revision.
IIS computation (Chinneck, 2008) identifies which constraints conflict but not why: an IIS containing an inventory balance equation and a capacity constraint at echelon 3 reveals a mathematical conflict, yet the analyst must still determine whether capacity was understated, demand was overstated, or lead time was misspecified—each calling for a different repair. Even a correct diagnosis does not guarantee a good repair: relaxing a capacity constraint may restore feasibility but let echelon 3 order ten times its physical throughput—a mathematically valid yet operationally meaningless fix. Unlike general software debugging, where unit tests sample program behavior, diagnosing and repairing supply chain models demands both mathematical reasoning (interpreting IIS output) and operational reasoning (judging whether the repaired solution makes physical sense).
A feasible, optimal solution is necessary but not sufficient. A repaired supply chain model must also satisfy what we call operational rationality: its solution should exhibit properties consistent with supply chain theory—base-stock structure (Clark and Scarf, 1960), bounded bullwhip amplification (Lee et al., 1997; Chen et al., 2000), inventory allocation, cost consistency (Zipkin, 2000), and order smoothing. These five checks form a RationalityOracle that filters out “technically correct but operationally absurd” repairs—a type of error that pure feasibility metrics miss entirely.
LLMs now address operations management tasks (Dai and Swaminathan, 2025; Chen et al., 2025; Zhao et al., 2025); benchmarks such as NL4Opt (Ramamonjison et al., 2023) and OptiBench (Yang et al., 2024) evaluate one-shot formulation, while PILOT-Bench (Ao et al., 2026c) adds multi-step tool interaction but targets formulation, not repair of faulty models. The iterative diagnosis-and-repair process that consumes analyst time in practice remains unevaluated. LLMs exhibit a 64.5% self-correction blind spot rate (Tsui, 2025). Existing correction benchmarks either exclude OR entirely (Tie et al., 2025) or rely on unit tests rather than deterministic solver feedback (Jimenez et al., 2024). Can LLMs diagnose and repair supply chain optimization models effectively when given structured solver feedback?
A closed-loop framework linking optimization feasibility to operational decisions.
OptiRepair splits optimization model diagnosis and repair into two independently deployable phases. Phase 1 iterates on solver IIS feedback to diagnose and repair constraint conflicts (up to 20 turns); it applies to any LP, not just supply chains. Phase 2 applies a domain oracle that checks whether the solution is operationally sound; we implement it for supply chains with five theory-grounded rationality checks (Section 3.3.2). Each phase is independently trainable, but combined they connect optimization (“is the model solvable?”) to operations (“does the solution make sense?”). Two evaluation tracks test the framework: OptiRepair-SC runs the full two-phase pipeline on 976 supply chain problems (10 error types, ME-1ME-10); OptiRepair-Debug isolates Phase 1 on $\sim$}5,000 generic LP problems (9 error types, four OR domains). This work extends a solver-in-the-loop evaluation (Ao et al., 2026d) with supply chain specialization, two-phase rationality validation, and a training pipeline benchmarked against 22 frontier API models (Table 2).
A domain-specialized training pipeline that outperforms frontier API models.
Using iterative self-taught reasoning (STaR) with solver-verified rewards, we train two 8B-parameter models—one for feasibility repair (Phase 1) and one for rationality repair (Phase 2). The combined pipeline reaches 81.7% Rational Recovery Rate (RRR)—the fraction of problems where the agent restores solver feasibility and produces an operationally rational solution. The best API model (GPT-5.2, tied with Gemini 2.5 Pro) reaches 42.2%, a gap of 39.5 percentage points. The gap holds across two base architectures: Qwen3-8B at 81.7% and Llama-3.1-8B at 80.3%, differing by only 1.4 percentage points.
Empirical evidence that domain specialization, not model scale, drives operational value.
Across 22 API models spanning 7 families, two gaps persist: solver interaction (API average 27.6% recovery rate versus 97.2% for trained models) and operational rationale (roughly one in four feasible repairs violate supply chain theory); prompt engineering cannot reliably close either (one model improves by 7.4pp, another degrades by 24.1pp). Domain-specialized 8B models outperform all 22 frontier API models; codifying operational rationale as solver-verifiable checks improved results more than scaling model parameters.
Table 1 summarizes the central result. Section 2 reviews related work; Section 3 defines the evaluation framework; Section 4 details the training pipeline; Section 5 reports results; and Section 6 discusses implications and limitations.
Table 1: Central result on OptiRepair-SC (284-problem test set). RRR: fraction of problems where the model restores feasibility and produces an operationally rational solution.
Configuration RRR
Trained 8B pipeline (best) 81.7%
Best API (GPT-5.2 / Gemini 2.5 Pro) 42.2%
Average API (22 models, 7 families) 21.3%
2 Related Work
OptiRepair draws on three research streams: AI and LLMs in operations management, LLM-assisted optimization, and supply chain theory together with verifiable-reward training methods.
2.1 AI and LLMs in Operations Management
Dai and Swaminathan (2025) organize AI in OM around three pillars: AI for OM (using AI to improve operations), OM for AI (applying operational principles to AI system design, e.g., Ao et al., 2025), and Human-AI interaction (how humans and AI collaborate on operational tasks). OptiRepair belongs to the first pillar—training AI agents for an operational task—and advances the third by identifying where AI succeeds (Phase 2 rationality repair, 76.3% average pass rate) and where it fails (Phase 1 repair, 27.6% average recovery rate).
Behavioral operations research identifies systematic biases in human operational decisions. Schweitzer and Cachon (2000) established that newsvendor ordering exhibits a pull-to-center bias, and Bolton and Katok (2008) confirmed that experience reduces but does not eliminate these biases. Demand amplification studies (Sterman, 1989) and behavioral operations surveys (Donohue et al., 2020) reinforce the case for automated diagnosis assistance. Several studies now test whether LLMs replicate these biases. Chen et al. (2025) reported that ChatGPT exhibits demand-chasing and pull-to-center biases in newsvendor decisions, replicating patterns Schweitzer and Cachon (2000) and Kremer et al. (2010) observed in human subjects. AIM-Bench (Zhao et al., 2025) tested agentic LLMs as inventory managers in multi-period settings, confirming persistent pull-to-center behavior. Zhang et al. (2025) evaluated LLMs as human-behavior simulators in OM experiments, finding that they predict treatment effects but miss distributional properties of actual decisions. In concurrent work, Baek et al. (2026) benchmarked LLM agents on over 1,000 inventory scenarios, finding that OR-augmented LLM policies and human-AI teams each outperform their components in isolation; Simchi-Levi et al. (2026) showed that LLMs can serve as interactive interfaces to supply chain solvers, explaining recommendations and adjusting parameters through dialogue; and Long et al. (2025) deployed LLM agents on the beer distribution game, observing severe bullwhip effects in decentralized settings. These studies confirm that LLMs and optimization tools are complementary—a premise OptiRepair applies through solver-in-the-loop repair. Gijsbrechts et al. (2022) applied deep RL to inventory management across dual sourcing, lost sales, and multi-echelon settings, finding that RL policies match or exceed heuristics but require careful reward design. Davis et al. (2024) argued for integrating machine learning and behavioral science, since neither approach alone captures how humans make operational decisions. Di Sorbo et al. (2025) measured how warnings and endorsements shape human-AI collaboration in demand forecasting. More broadly, integrating ML-generated signals into sequential decisions under imperfect feedback—whether LLM judges replacing human evaluators, prediction-powered optimization, or surrogate rewards for bandits—is an active area (Ao et al., 2026a, b; Ji et al., 2025); our RationalityOracle takes a similar approach, replacing expert review with automated solver-grounded checks. OptiRepair differs from these lines of work in two ways: it evaluates AI agents on a structured repair task with deterministic solver verification rather than prediction or quantity decisions, and it treats operational rationality as a first-class metric separate from task completion.
2.2 LLM-Assisted Optimization
Multiple benchmarks now evaluate LLMs on optimization tasks. Most focus on one-shot formulation: NL4Opt (Ramamonjison et al., 2023), OptiBench (Yang et al., 2024), ORLM (Huang et al., 2025), MAMO (Huang et al., 2024), and others (Xiao et al., 2024; AhmadiTeshnizi et al., 2024). Several testbeds add multi-step interaction: PILOT-Bench (Ao et al., 2026c) evaluates LLM robustness under probabilistic tool failures and variable instruction quality, DP-Bench (Zhou et al., 2025) applies solver verification to dynamic programming formulation, and LEAN-LLM-OPT (Liang et al., 2026) coordinates multiple agents to auto-formulate large-scale problems via structured workflow; all target formulation accuracy rather than repairing faulty models. Tsui (2025) measured a 64.5% self-correction blind spot rate in LLMs; CorrectBench (Tie et al., 2025) benchmarks correction across reasoning domains but excludes operations research. SWE-bench (Jimenez et al., 2024) benchmarks iterative code debugging, but unit tests only sample program behavior, whereas a solver delivers complete, deterministic feedback. OptiRepair addresses this gap: it evaluates iterative diagnosis and repair of faulty models with solver feedback, extending the correctness criterion beyond feasibility to domain-specific operational rationality.
Table 2: Evaluation framework comparison. OptiRepair-SC is the first to evaluate both iterative diagnosis-and-repair and operational rationality in a closed-loop supply chain setting.
Framework Task Oracle Multi-step Self-Corr. Rationality
[…]
OptiBench (Yang et al., 2024) Formulation
MAMO (Huang et al., 2024) Complex LP
ORLM (Huang et al., 2025) Formulation
LEAN-LLM-OPT (Liang et al., 2026) Formulation Solver
AIM-Bench (Zhao et al., 2025) Inventory Closed-form
InventoryBench (Baek et al., 2026) Inventory Closed-form
PILOT-Bench (Ao et al., 2026c) Tool Workflow Execution
SWE-bench (Jimenez et al., 2024) Code Debug Unit Tests Limited
CorrectBench (Tie et al., 2025) Self-Corr. General
DP-Bench (Zhou et al., 2025) DP Problems Solver
OptiRepair-Debug OR Repair Solver IIS
OptiRepair-SC SC Repair Solver IIS + Oracle
Table 2 contrasts OptiRepair with existing benchmarks on oracle type, multi-step interaction, self-correction, and domain-specific rationality checks.
2.3 Supply Chain Theory and Training Methods
Our error classification and rationality checks build on classical supply chain theory. Clark and Scarf (1960) proved that echelon base-stock policies optimize serial multi-echelon systems. Lee et al. (1997) traced order variance amplification to information distortion, and Chen et al. (2000) quantified the resulting bullwhip effect. Zipkin (2000) consolidated the foundations of inventory management. Our ten error types and five rationality checks encode these theoretical properties; Section 3.2 gives the detailed mapping.
Our training approach extends reinforcement learning with verifiable rewards (RLVR). DeepSeek-AI (2025) showed that GRPO with outcome verification induces reasoning behavior without step-level supervision; Shao et al. (2024) and Lambert et al. (2025) confirmed that RLVR generalizes across reasoning domains. Zelikman et al. (2022) introduced STaR for bootstrapped self-improvement. We adapt these methods to OR model repair, where Gurobi supplies exact, deterministic verification at every step—feasibility status, objective value, and IIS certificates—enabling fully automated reward computation without human annotation.
3 Closed-Loop Diagnosis and Repair Framework
OptiRepair decomposes optimization model repair into a domain-agnostic diagnosis-and-repair phase and a domain-specific validation phase. Phase I applies to any LP: given an infeasible model and solver IIS feedback, the agent iteratively diagnoses constraint conflicts and repairs the formulation. Phase II invokes a domain oracle that checks whether the repaired solution satisfies operational criteria; for supply chain models, the oracle verifies properties grounded in inventory theory (Clark and Scarf, 1960; Chen et al., 2000).
3.1 Multi-Echelon LP Formulation
Consider a serial supply chain with $N$ echelons indexed $n=1,\ldots,N$, where echelon 1 is the retailer (most downstream) and echelon $N$ is the factory (most upstream). Material flows from echelon $N$ toward echelon 1; orders flow in the opposite direction. The planning horizon spans $T$ discrete periods.
Each echelon $n$ incurs holding cost $h_{n}$ and backorder cost $b_{n}$ per unit per period, faces production or order capacity $C_{n}$, and has replenishment lead time $L_{n}$ periods. Customer demand at the retailer in period $t$ is $d_{t}$. Decision variables are order quantity $x_{n,t}$, on-hand inventory $I_{n,t}$, and backorders $B_{n,t}$ at each echelon and period. The LP minimizes total holding and backorder costs:
minimize $\sum_{n=1}^{N}\sum_{t=1}^{T}\bigl(h_{n}\,I_{n,t}+b_{n}\,B_{n,t}\bigr)$ (1)
s.t. $I_{n,t}-B_{n,t}=I_{n,t-1}-B_{n,t-1}+x_{n,t-L_{n}}-d_{n,t},\quad\forall\,n,\,t$ (2)
$x_{n,t}\leq C_{n},\quad\forall\,n,\,t$ (3)
subject also to demand propagation ($d_{n,t}=x_{n-1,t}$ for $n\geq 2$; $d_{1,t}=d_{t}$) and non-negativity constraints ($x_{n,t},\,I_{n,t},\,B_{n,t}\geq 0$); see Appendix A for the full specification. Constraint Eq. (2) enforces flow conservation at each echelon, while demand propagation transmits order variability upstream—the bullwhip effect (Lee et al., 1997; Chen et al., 2000; Simchi-Levi et al., 2008). Appendix A gives the complete formulation with all constraint families.
Our generator produces instances with 2–5 echelons, 12–24 periods, and varying cost structures and demand patterns (stationary, step-change, and seasonal). Typical instances contain 70–400 decision variables and 60–350 constraints. Gurobi solves every instance to optimality before error injection, confirming a feasible baseline.
3.2 Error Classification
Table 3: Supply chain error classification. Ten error types (ME-1ME-10) represent common modeling mistakes in multi-echelon inventory planning.
Type Error Name Operational Meaning Difficulty
ME-1 Demand Inflation Overstated downstream demand Hard
ME-2 Lead Time Error Misspecified replenishment lead time Hard
ME-3 Balance Violation Impossible inventory balance state Easy
ME-4 Capacity Reduction Production capacity below demand Hard
ME-5 Cost Structure Error Non-monotonic holding costs Easy
ME-6 Bullwhip Amplification Forced upstream order amplification Medium
ME-7 Coefficient Perturbation Distorted arrival coefficient Medium
ME-8 Sign Error Inverted demand propagation sign Medium
ME-9 Redundant Constraint Conflicting minimum-order bound Medium
ME-10 Index Mismatch Wrong time index in demand propagation Hard
Table 3 defines ten error types (ME-1ME-10) covering common modeling mistakes in multi-echelon inventory planning (Simchi-Levi et al., 2008; Zipkin, 2000). Each type targets a specific LP component; we group them into five categories by the operational subsystem they affect.
Demand and timing errors (ME-1, ME-2) produce multi-echelon IIS patterns requiring upstream–downstream dependency tracing. Balance and capacity errors (ME-3, ME-4): ME-3 forces an impossible inventory balance state (localized, easy-to-identify IIS signal), while ME-4 sets capacity far below demand, producing cascading infeasibility across periods. Cost and structural errors (ME-5, ME-6): ME-5 distorts holding costs, yielding feasible but operationally irrational solutions; ME-6 adds bullwhip-inducing constraints that cascade through the echelon structure. Coefficient and sign errors (ME-7, ME-8) perturb arrival coefficients or invert demand propagation signs, creating multi-constraint IIS patterns. Constraint and index errors (ME-9, ME-10): ME-9 adds redundant order constraints; ME-10 corrupts time indices in demand propagation, producing misleading IIS signals.
A saboteur module generates all errors programmatically: given a valid LP instance, it selects an error type, applies a controlled perturbation, and verifies that (i) the sabotaged model returns Infeasible (or Optimal-but-irrational for ME-5), (ii) the IIS is non-empty, and (iii) a ground-truth fix restores both feasibility and rationality.
3.3 Two-Phase Closed-Loop MDP
OptiRepair uses a two-phase closed-loop architecture (Figure 1). Phase I is domain-agnostic: the same state space, action space, and transitions apply to any infeasible LP regardless of domain. Phase II applies a domain-specific oracle that checks operational correctness. Each phase operates as a self-contained decision loop with its own reward signal; combined, they connect optimization (“is this model solvable?”) to operations (“does this solution support sound decisions?”). We model both phases as a deterministic Markov Decision Process $(\mathcal{S},\mathcal{A},T,R,\gamma)$.
Figure 1: Two-phase closed-loop architecture of OptiRepair. Phase I (domain-agnostic) iteratively diagnoses and repairs an infeasible LP via solver IIS feedback. Upon reaching Optimal, Phase II (domain-specific) validates operational correctness through a replaceable domain oracle—implemented here with five supply chain rationality checks.
3.3.1 Phase I: Diagnose and Repair.
The agent receives an infeasible LP and a natural-language description of the intended model. At each step, the agent observes the state, selects an action, and the environment executes it via Gurobi.
State space.
A state $s\in\mathcal{S}$ is a tuple:
$s=(\text{problem}_{\text{NL}},\;\text{code},\;\text{status},\;\text{iis},\;\text{history},\;t),$ (4)
where $\text{problem}_{\text{NL}}$ is the natural-language problem description, code is the current Gurobi model code, $\text{status}\in\{\textsc{Infeasible},\textsc{Optimal},\ldots\}$ is the solver return status, iis lists the constraints and variable bounds in the irreducible infeasible subsystem, history records previous actions and outcomes, and $t$ counts steps.
Action space.
Actions $a\in\mathcal{A}$ split into three categories: diagnostic actions (GET_IIS, CHECK_SLACK) that query the solver without modifying the model; repair actions (RELAX_CONSTRAINT, DROP_CONSTRAINT, UPDATE_OBJ, UPDATE_BOUNDS, UPDATE_RHS) that modify the model; and the meta action SUBMIT that terminates the episode. Each action specifies a target constraint or variable by name and, where applicable, a numerical value.
Transition.
Transitions are deterministic: given state $s$ and action $a$, the environment applies $a$ to the Gurobi model, re-solves, recomputes the IIS if still infeasible, and returns $s^{\prime}$. Phase I succeeds when the solver returns Optimal. Exhausting the 20-step budget without restoring feasibility ends the episode as a Phase I failure. Figure 2 illustrates a Phase I interaction: the agent diagnoses an ME-4 error (capacity too tight) using GET_IIS and repairs it in a single step.
Figure 2: Phase I repair interaction. The agent receives an infeasible supply chain LP (Gurobi), uses GET_IIS to diagnose conflicting constraints, and repairs the capacity bound. When the solver returns Optimal, the episode transitions to Phase II. Full system prompt and transcript appear in Appendix D.
3.3.2 Phase II: Validate.
Once Phase I reaches Optimal, the repaired model enters Phase II. A replaceable domain oracle encodes expert criteria for operational correctness beyond mathematical feasibility. For multi-echelon supply chain models, our RationalityOracle enforces five checks grounded in inventory theory (Clark and Scarf, 1960; Chen et al., 2000): (1) base-stock structure (inventory levels approximate echelon base-stock policies), (2) bullwhip ratio (order variance amplification stays within theoretical bounds (Dejonckheere et al., 2003)), (3) inventory allocation (downstream echelons do not hoard disproportionate inventory), (4) cost consistency (holding cost ordering and objective coefficients match the problem specification), and (5) order smoothing (order quantities avoid extreme oscillation). Appendix C gives formal definitions and thresholds.
Not all checks apply to every error type: an error-type-aware mapping excludes checks whose violations reflect saboteur-injection artifacts rather than genuine irrationality (Appendix C gives the full mapping). When any applicable check fails, the agent receives natural-language feedback identifying the violated property and re-enters the repair loop for up to three iterations. Phase II succeeds when all applicable checks pass; it fails if iterations run out or a repair breaks feasibility. Figure 3 shows a Phase II loop-back: the oracle detects a cost-consistency violation, the agent corrects the objective coefficient, and the oracle confirms all checks pass.
Figure 3: Phase II rationality repair interaction. The rationality oracle detects that the Phase I repair created a cost-consistency violation: the high-cost retailer holds more inventory than the cheaper warehouse. The agent corrects the objective coefficient; the solver re-optimizes and the oracle confirms all five checks pass.
3.3.3 Combined reward.
The episode reward $R=R_{\text{outcome}}+R_{\text{rationality}}$ sums a feasibility term (+100+100 if Optimal, 50-50 otherwise) and a rationality term (+50+50 if all checks pass, 25-25 if Optimal but checks fail, 0 if Infeasible):
[…]
The agent therefore prioritizes feasibility restoration while still earning credit for solution quality.
3.4 Evaluation Metrics
Three complementary metrics measure repair performance.
Recovery Rate (RR).
The fraction of problems where the agent restores the solver to Optimal status within the step budget:
$\text{RR}=\frac{|\{i:\text{status}_{i}=\textsc{Optimal}\}|}{N}.$ (6)
RR isolates raw repair ability: can the agent find any feasible modification, regardless of solution quality?
Rational Recovery Rate (RRR).
The fraction of problems where the agent achieves Optimal and passes all applicable rationality checks:
$\text{RRR}=\frac{|\{i:\text{status}_{i}=\textsc{Optimal}\;\land\;\text{rational}_{i}=\text{True}\}|}{N}.$ (7)
RRR is the primary metric for OptiRepair-SC, encoding both repair skill (Phase I) and operational understanding (Phase II) in a single number.
Phase 2 Pass Rate (P2Pass).
Among problems where the agent achieves Optimal, the fraction that also pass rationality: $\text{P2Pass}=\text{RRR}/\text{RR}$. A model with high RR but low P2Pass restores feasibility yet yields operationally nonsensical solutions. Decomposing $\text{RRR}=\text{RR}\times\text{P2Pass}$ separates repair skill from operational understanding; Section 5 reports all three metrics.
3.5 Problem Set Construction
[…]
OptiRepair-SC.
OptiRepair-SC (OptiRepair Supply Chain) contains 976 instances across all ten error types (ME-1ME-10), with disjoint supply chain configurations between the training and test splits (Appendix G details the split). Each problem bundles a natural-language description, Gurobi Python code for the sabotaged LP, the ground-truth fix, and the error type label. Instances span 2–5 echelons, 12–24 periods, three demand patterns, holding-to-backorder ratios from 1:21{:}2 to 1:101{:}10, and both capacitated and uncapacitated variants. We validate every problem: the sabotaged code returns Infeasible (or Optimal-but-irrational for ME-5), and the ground-truth fix passes all applicable rationality checks.
OptiRepair-Debug.
To test whether Phase I transfers beyond supply chains, we include OptiRepair-Debug, which contains approximately 5,000 problems across nine generic LP error types (Types A–I) drawn from facility location, transportation, scheduling, and network flow models. OptiRepair-Debug exercises Phase I only—no Phase II oracle—confirming that the repair loop generalizes across domains without supply-chain-specific modifications. Section 5.2 summarizes OptiRepair-Debug results.
4 Training Pipeline
Two specialized 8B-parameter models drive the OptiRepair closed-loop pipeline—one per phase. The combined training procedure, OptiSTaR (Optimization Self-Taught Reasoning), cycles through STaR exploration seeded by solver feedback and GRPO refinement at each iteration. Section 4.1 motivates the two-model split; Section 4.2 details the iterative STaR procedure for Phase I repair; Section 4.3 covers oracle-based Phase II training; and Section 4.4 tests architecture generalization.
4.1 Two-Model Architecture
The OptiRepair pipeline uses two independently trained 8B-parameter models: a P1 model for Phase I repair (Infeasible $\to$ solver-feasible) and a P2 model for Phase II rationality repair (solver-feasible $\to$ operationally rational). Each model solves a self-contained problem with its own reward signal: P1 maximizes solver feasibility recovery using IIS feedback, while P2 maximizes oracle check compliance using domain-specific rationality criteria. Separating the two objectives isolates feasibility restoration from operational validation. Phase I demands LP structure and IIS-based diagnosis—skills that transfer across optimization domains (OptiRepair-Debug confirms this on four problem families). Phase II demands domain-specific knowledge (base-stock structure, bullwhip bounds, cost consistency) and can be swapped or extended without retraining Phase I. A joint model mixes both objectives: it can restore feasibility by deleting cost constraints, producing irrational solutions.
Both models initialize from Qwen3-8B-Instruct, selected via a pilot study (+41.9% $\text{RR@}{5}$ headroom). At inference, P1 runs for up to 20 turns; once the solver returns Optimal, P2 takes over for up to 3 rationality repair iterations.
4.2 Iterative STaR for Phase I Repair
Figure 5: OptiSTaR training pipeline. Top: Phase I iterates through beam search exploration (K=32K\!=\!32 candidates), STaR distillation, and GRPO refinement; three iterations raise $\text{RR@}{5}$ from 44.4% to 75.7%. Gurobi provides IIS feedback for beam search and outcome rewards for GRPO. Middle right: Phase II trains independently on oracle-generated data (SFT) refined by solver-based GRPO. Bottom: The two independently trained 8B models combine into the final pipeline ($\text{RRR}=81.7\%$).
Phase I training uses an iterative self-taught reasoning (STaR) loop (Zelikman et al., 2022), cycling through beam search exploration (Wei et al., 2022), distillation, and reinforcement learning in three stages (Figure 5).
Stage 1: Beam Search Exploration.
The current model generates $K=32$ candidate trajectories per problem; the shortest trajectory that restores Optimal status with $\text{DA}\geq 0.5$ survives.
Stage 2: Supervised Distillation.
SFT on successful beam trajectories with completion-only loss distills search-time gains into the model’s greedy policy.
Stage 3: GRPO Refinement.
Group Relative Policy Optimization (Shao et al., 2024; DeepSeek-AI, 2025) refines the distilled model. For each problem, the model generates $G=8$ rollout trajectories; rewards normalize within each group to yield advantages. The composite reward balances three objectives:
$R=0.5\cdot R_{\text{outcome}}+0.3\cdot R_{\text{diagnosis}}+0.2\cdot R_{\text{efficiency}},$ (8)
where $R_{\text{outcome}}=+100$ if the solver returns Optimal and 50-50 otherwise; $R_{\text{diagnosis}}=\text{DA}\times 100$, rewarding overlap between the model’s targeted constraints and the ground-truth IIS; and $R_{\text{efficiency}}=-1$ per repair step, penalizing long trajectories. A faithfulness penalty of 20-20 fires when the model targets constraints outside the IIS, discouraging indirect fixes that mask the root cause.
Hyperparameters: KL coefficient $\beta=0.02$, asymmetric clipping $\epsilon_{\text{low}},\epsilon_{\text{high}}]=[0.2,0.28]$, LoRA rank $r=16$, $\alpha=32$.
Iterative Loop.
The STaR loop ran three iterations, seeded with 696 successful trajectories from GPT-5.2, o4-mini, and DeepSeek-R1 teachers. Beam trajectories for hard error types (ME-2, ME-4) are filtered against ground-truth rationality: trajectories whose repairs fail rationality checks are replaced with upsampled ground-truth traces.
Training Progression.
Three STaR iterations raise Phase I $\text{RR@}{5}$ from 44.4% to 75.7% and P1 RR to 97.2%.
The Indirect Role of Reinforcement Learning.
GRPO’s direct effect on greedy performance is marginal (+0.7pp in $\text{RR@}{5}$). Its value is indirect: GRPO refinement raises beam search solve rate from 81.4% to 98.3%, feeding higher-quality distillation data into the next STaR iteration. In multi-turn agentic settings, RL amplifies data quality rather than directly improving the greedy policy.
4.3 Oracle-Based Phase II Training
Phase II training teaches the model to recognize and repair operationally irrational solutions that remain mathematically feasible. The critical design choice is the rationality oracle—the automated verifier that defines “correct” Phase II behavior.
Oracle Design Determines Effectiveness.
Three oracle design decisions—error-type-aware check activation, artifact-check relaxation, and prefix-matched constraint operations—together determine Phase II effectiveness. Precise correctness rules are a prerequisite for effective training.
Solver-Based GRPO Reward.
Phase II GRPO grounds its reward in the solver: each rollout trajectory runs in Gurobi and the RationalityOracle scores it, tying the training signal to actual operational quality rather than textual similarity. The solver-based reward uses LoRA $\alpha=0.5$ because the signal is binary: the solution either passes rationality checks or fails. Paired with GRPO-trained Phase 1, the final pipeline reaches $\text{RRR}=81.7\%$ on the held-out test set.
Error-Type Specialization.
ME-5 errors (pure Phase II problems) reach 100% RRR. The hardest errors, ME-2 and ME-4, require both phases and yield the largest GRPO gains: from 6% and 27% RRR to 64.5% and 61.3%, respectively (Section 5.4).
4.4 Architecture Generalization
Replicating the full pipeline on Llama-3.1-8B-Instruct yields $\text{RRR}=80.3\%$ on the test set—within 1.4 percentage points of Qwen3-8B’s 81.7%, with identical Phase I recovery rates (97.2%). The ablation pattern also holds: replacing trained P1 with Claude Opus 4.5 drops RRR from 80.3% to 40.5%. Full results appear in Section 5.6. Because the two architectures differ in tokenizer, pre-training data, and instruction-tuning methodology, this consistency confirms that the gains stem from the training procedure and oracle specification rather than architecture-specific properties.
5 Experiments
Evaluation spans two tracks: OptiRepair-Debug (general OR repair, single-phase) and OptiRepair-SC (supply chain closed-loop, two-phase).
5.1 Experimental Setup
Evaluation tracks. OptiRepair-Debug tests single-phase repair on 450 stratified samples per model across nine error types (Types A–I), drawn from linear and integer programs spanning production planning, transportation, scheduling, and facility location. OptiRepair-SC tests the full two-phase closed-loop pipeline on 284 test problems (from the 976-problem set) across ten supply chain error types (ME-1ME-10), with a disjoint training split for domain-specialized models.
Models. The evaluation covers 30 models. OptiRepair-Debug tests 4 locally trained Qwen3-8B variants (SFT, OptiSTaR, DAPO, Curriculum) and 22 API models from seven families: OpenAI (7), Google (3), Anthropic Claude (4), DeepSeek (2), Qwen (4), Meta (1), and Moonshot (1). OptiRepair-SC tests the same 22 API models plus two trained 8B-parameter pipelines (Qwen3-8B and Llama-3.1-8B).
[…]
5.2 OptiRepair-Debug Results: General OR Repair
OptiRepair-Debug tests whether Phase I’s domain-agnostic repair loop transfers beyond supply chains, spanning approximately 5,000 problems across 9 generic LP error types and 26 models. The trained Qwen3-8B-OptiSTaR reaches 95.3% RR@5 and 62.4% DA, exceeding the best API model (o4-mini at 86.2% RR@5) by 9.1 percentage points. Trained models also complete repairs in 2.25 steps on average versus 3.78 for the best API models. Full per-model and per-error-type breakdowns are available from the authors.
5.3 OptiRepair-SC Results: Supply Chain Closed-Loop
Table 4: OptiRepair-SC results: 22 API models on the supply chain closed-loop pipeline (284 test problems). Models ranked by Rational Recovery Rate (RRR).
# Model $R$ RRR P2Pass Steps # Model $R$ RRR P2Pass Steps
1 GPT-5.2 53.5 42.2 79.0 14.0 12 Claude Haiku 4.5 20.8 19.0 91.5 7.4
2 Gemini 2.5 Pro 53.9 42.2 78.4 9.3 13 GPT-o3 19.7 19.0 96.4 7.5
3 DeepSeek-R1-0528 53.2 37.0 69.5 13.8 14 GPT-4.1 18.0 17.6 98.0 5.9
4 Gemini 2.5 Flash 45.1 32.0 71.1 9.5 15 GPT-4.1 Mini 17.2 17.2 100.0 6.6
5 Qwen2.5-7B 41.2 31.0 75.2 8.1 16 GPT-5 Mini 18.7 16.9 90.6 6.8
6 Claude Opus 4.5 33.5 30.3 90.5 16.6 17 GPT-o1 14.9 13.5 90.5 4.4
7 Qwen2.5-Max 32.8 26.8 81.7 7.0 18 Claude 3.7 Sonnet 13.7 13.4 97.4 4.5
8 GPT-o4-mini 35.9 25.7 71.6 9.5 19 Qwen2.5-14B 13.0 12.7 97.3 3.7
9 Claude Sonnet 4 26.4 23.6 89.3 7.2 20 Qwen2.5-32B 23.2 4.2 18.2 5.5
10 DeepSeek-V3.2 23.9 21.8 91.2 17.0 21 Kimi K2 15.1 1.4 9.3 5.3
11 Gemini 2.0 Flash 22.9 20.1 87.7 5.9 22 Llama-3.3-70B 11.3 0.4 3.1 2.1
Average (22 models) 27.6 21.3 76.3 7.8
[…]
Model Arch. P1 RR RRR ME-2 ME-4 ME-5 Steps Tokens
Qwen3-8B (OptiSTaR) Qwen 97.2 81.7 64.5 61.3 100.0 5.2 3,439
Llama-3.1-8B (OptiSTaR) Llama 97.2 80.3 67.7 74.2 100.0 4.3 3,180
Best API (GPT-5.2) 53.5 42.2 41.9 38.7 100.0 14.0 20,730
Avg. API (22 models) 27.6 21.3 3.7 4.7 87.3 7.8 30,078
Figure 6: Rational Recovery Rate (RRR) for 22 API models and two trained 8B pipelines (Qwen3-8B and Llama-3.1-8B) on OptiRepair-SC (284-problem test set). The vertical dashed line marks the API average (21.3%). Both trained models outperform the best API by $\geq$38 percentage points.
The trained Qwen3-8B pipeline reaches 81.7% RRR (Table 5), beating the best API models—GPT-5.2 and Gemini 2.5 Pro, both at 42.2% (Table 4)—by 39.5 percentage points; the average API model manages only 21.3%. Performance varies widely across the 22 API models (Figure 6). At the bottom, Qwen2.5-32B (4.2% RRR), Kimi K2 (1.4%), and Llama-3.3-70B (0.4%) fall below 5% due to Phase 2 format compliance failures. GPT-5.2, Gemini 2.5 Pro, and DeepSeek-R1-0528 all exceed 37% RRR, each with Phase 1 recovery above 53%. Ranks 4–12 cluster between 19% and 32%; model family and parameter count are weak predictors of performance.
The gap concentrates in Phase 1: API models average 27.6% RR versus 97.2% for trained models—a 69.6 percentage point gap (Figure 7). On OptiRepair-Debug, the best API models exceed 86% RR@5 on general OR problems; the drop to 28% on supply chain problems reflects harder repair challenges: longer constraint chains, multiple interacting echelons (Simchi-Levi et al., 2008), and domain-specific variable naming. Once feasibility is restored, the validation phase poses a lower barrier: the 19 format-compliant API models average 86.7% P2Pass, with several above 90%. The bottleneck is finding a feasible repair, not judging its operational quality.
Figure 7: Phase 1 Recovery Rate vs. Phase 2 Pass Rate for all 22 API models and two trained 8B pipelines. Bubble size $\propto$ RRR. API models cluster upper-left (high P2Pass, low $R$); trained models reach the upper-right corner. The vertical gap between clusters—roughly 70 percentage points in RR—confirms that Phase 1 repair drives the overall performance difference.
Model families exhibit distinct repair profiles. Claude models achieve the highest P2Pass (89–97%) but the lowest RR among top-tier models (13–34%); GPT-5.2 and Gemini 2.5 Pro balance both (RR >>53%, P2Pass $\approx$79%). Reasoning-specialized models (o3: 19.0%, o1: 13.5%) do not outperform general models. DeepSeek-R1-0528 ranks third (37.0%) with strong RR (53.2%) but lower P2Pass (69.5%), illustrating a feasibility–rationality trade-off: aggressive repairs restore solvability at the cost of operational quality.
5.4 Per-Error-Type Analysis
Table 9 (Appendix E) breaks down RRR by error type for the top-6 API models and both trained models. The ten error types range from universally solvable to near-impossible for API models, with a sharp boundary between errors that produce direct diagnostic signals and those requiring multi-echelon causal reasoning.
All top models score 90–100% RRR on ME-5 (cost structure error): the model is already feasible, and the rationality oracle directly identifies the non-monotonic holding costs, making the repair unambiguous.
ME-2 (lead time error) and ME-4 (capacity reduction) separate trained from API models. On ME-2, the best API model (GPT-5.2) reaches 41.9% RRR vs. trained Qwen3-8B at 64.5%; on ME-4, GPT-5.2 manages 38.7% vs. Llama-3.1-8B at 74.2%. These errors require tracing how lead-time coupling and capacity constraints interact with inventory balance equations across echelons—domain knowledge that general models lack.
ME-3 (balance violation) tops the API success ranking (63–77% for top models) because the IIS directly pinpoints the violated constraint and the repair is localized. ME-1 (demand inflation) and ME-10 (index mismatch) hover near zero for API models (0–7%), as these errors generate misleading IIS signals that require backward reasoning through the echelon structure. The pattern is clear: when the IIS points directly to the root cause (ME-3, ME-5), general models suffice; when diagnosis requires tracing interactions across echelons (ME-1, ME-2, ME-4, ME-10), only specialized training closes the gap.
5.5 Component Ablation
Table 6: Component ablation on OptiRepair-SC (284 test problems). Phase 1 repair is the critical bottleneck: replacing the trained P1 model with API alternatives reduces RRR by 22–40 percentage points. The pattern holds across both architectures.
P1 Model P2 Model P1 RR (%) RRR (%)
Qwen3-8B pipeline (8B)
Qwen3-8B OptiSTaR Qwen3-8B P2 OptiSTaR 97.2 81.7
Qwen3-8B OptiSTaR Claude Opus 4.5 97.2 69.4
Qwen3-8B OptiSTaR GPT-5.2 96.8 68.7
Claude Opus 4.5 Qwen3-8B P2 OptiSTaR 55.6 45.1
GPT-5.2 Qwen3-8B P2 OptiSTaR 63.0 55.6
Llama-3.1-8B pipeline (8B)
Llama-3.1-8B OptiSTaR Llama-3.1-8B P2 OptiSTaR 97.2 80.3
Llama-3.1-8B OptiSTaR Claude Opus 4.5 96.8 71.1
Llama-3.1-8B OptiSTaR GPT-5.2 97.2 71.1
Claude Opus 4.5 Llama-3.1-8B P2 OptiSTaR 51.1 40.5
GPT-5.2 Llama-3.1-8B P2 OptiSTaR 66.5 58.1
Full API (no training)
GPT-5.2 GPT-5.2 53.5 42.3
Claude Opus 4.5 Claude Opus 4.5 33.5 30.3
To isolate each phase’s contribution, we substitute trained models with API alternatives (Table 6). Replacing the trained Phase 1 model cuts RRR by 22–40pp; replacing only Phase 2 costs 9–13pp. The trained Phase 1 model reaches 97.2% RR, exceeding both API alternatives by 31–42pp. API models perform adequately on Phase 2 because operational reasoning aligns more closely with their general capabilities than multi-step repair does. End-to-end, GPT-5.2 scores 42.3% RRR and Opus 4.5 scores 30.3%—Phase 1 repair drives the gap. For practitioners building in-house repair tools, Phase 1 capability is the higher-return investment.
5.6 Architecture Generalization
To test architecture dependence, we replicate the full training pipeline on Llama-3.1-8B (Llama-3.1-8B-Instruct). Llama-3.1-8B scores 80.3% RRR versus Qwen3-8B’s 81.7%—a 1.4pp gap—with identical Phase 1 recovery rates (97.2%). The ablation patterns replicate across architectures: trained Phase 1 outperforms API Phase 1 by 30–46pp, trained Phase 2 outperforms API Phase 2 by 9–13pp, confirming that the training procedure, not architecture-specific properties, drives effectiveness.
5.7 Prompt Engineering Limitations
Two guided prompt variants augment the baseline API prompt: an anti-loop variant with explicit loop-avoidance instructions, and a domain-knowledge variant with supply chain structural hints (base-stock intuition, echelon coupling; full text in Appendix F). The domain-knowledge prompt raises Claude Opus 4.5 by 7.4 percentage points but degrades GPT-5.2 by 24.1 percentage points; the anti-loop prompt hurts both models. Prompt engineering alone cannot close the 38–40 percentage point gap to domain-specialized training.
5.8 Repair Efficiency
[…]
Figure 8: Repair efficiency: average steps vs. average tokens per problem for all 24 models. Bubble size $\propto$ RRR. Trained 8B models cluster in the bottom-left corner—fewer steps, fewer tokens, higher RRR—while the highest-RRR API models (GPT-5.2, DeepSeek-R1) consume the most resources.
6 Discussion and Conclusion
OptiRepair tests whether AI agents can diagnose and repair supply chain optimization models through a closed-loop framework evaluated on 22 API models and two trained 8B-parameter pipelines. Domain-specialized 8B models achieve 80–82% RRR, outperforming the best API model (42.2%) by 38–40pp across both architectures. The gap concentrates in Phase 1 repair (API average: 27.6% RR vs. 97.2% trained), while Phase 2 rationality pass rates remain high (76.3%). Precise correctness specification, not model capability, determines Phase 2 effectiveness.
6.1 Managerial Implications
The recovery bottleneck is solver interaction.
API models pass Phase 2 rationality checks 76.3% of the time—frontier LLMs already reason well about supply chain operations. Yet those same models restore only 27.6% of infeasible formulations, because iterating on solver feedback, diagnosing IIS conflicts, and proposing targeted constraint repairs is a different skill from answering questions about inventory theory. Adding supply chain knowledge to prompts does not help reliably: one model gains 7.4pp, another loses 24.1pp. For operations teams evaluating AI for model maintenance, the scarcest capability is solver interaction—and it responds to targeted training on a modest compute budget.
Feasibility is not rationality: the AI-human decision gap.
Restoring feasibility without ensuring operational rationality risks producing solutions that appear valid but encode flawed logic—potentially worse than leaving the model for human review. Roughly one in four API repairs that restore solver feasibility violate an operational property—inverted cost hierarchies, inventory hoarding at upstream echelons, or demand-amplifying orders—that no experienced analyst would accept. The gap between AI output and expert judgment is not a capability problem; it is a specification problem. The rationale behind a good supply chain solution—base-stock adherence, echelon coupling, cost consistency—is knowledge that experienced analysts apply implicitly. AI systems cannot learn it unless it is explicitly codified. Codifying that rationale as solver-verifiable checks—five predicates grounded in inventory theory (Clark and Scarf, 1960; Chen et al., 2000; Lee et al., 1997)—improved results more than training larger models. For organizations adopting AI in operational planning, formalizing what “rational” means in their context is the higher-return investment.
Separating repair from validation creates reusable components.
Because Phase 1 (feasibility repair) is domain-agnostic—confirmed across four OR problem families—extending to a new domain requires only a new Phase 2 oracle encoding that domain’s correctness criteria. Organizations can share one repair model across domains while maintaining domain-specific validation.
6.2 Limitations and Future Work
Several limitations qualify these findings. The benchmark uses synthetic errors with serial topology and deterministic demand; real-world practice involves error combinations, network topologies, stochastic demand, and multi-product settings. The RationalityOracle encodes operational criteria specific to multi-echelon inventory; organizations with different priorities would need to customize the checks. All evaluations use greedy decoding, and alternative inference strategies may yield different model rankings.
Expanding the error classification to mixed-integer programs and nonlinear formulations, incorporating real industrial repair traces, and evaluating multi-agent architectures where specialized agents collaborate with domain experts would each extend the evaluation.
For structured operational repair, domain expertise codified as a correctness oracle matters more than model scale—across all error types, architectures, and configurations tested.
Appendix A Full Multi-Echelon LP Formulation
This appendix details the complete linear programming formulation for the multi-echelon serial supply chain model used in the OptiRepair-SC evaluation. The formulation follows the echelon-stock framework of Clark and Scarf (1960) and its LP relaxation. Let $\mathcal{N}=\{1,\ldots,N\}$ denote echelons (1 = retailer, $N$ = factory; $N\in\{2,3,4,5\}$) and $\mathcal{T}=\{1,\ldots,T\}$ denote planning periods ($T\in\{12,16,20,24\}$).
A.1 Parameters
  • $h_{n}\geq 0$: holding cost per unit per period at echelon $n$. Range: $h_{n}\in[1,10]$.
  • $b_{n}\geq 0$: backorder cost per unit per period at echelon $n$. Range: $b_{n}\in[5,50]$.
  • $C_{n}>0$: production or order capacity at echelon $n$. Range: $C_{n}\in[50,500]$.
  • $L_{n}\geq 0$: replenishment lead time (in periods) at echelon $n$. Range: $L_{n}\in\{1,2,3\}$.
  • $d_{t}>0$: external customer demand at the retailer in period $t$.
  • $I_{n,0}^{\text{init}}\geq 0$: initial on-hand inventory at echelon $n$.
Demand patterns include three types: (i) stationary ($d_{t}=\bar{d}$ for all $t$), (ii) step-change ($d_{t}$ shifts from $\bar{d}_{1}$ to $\bar{d}_{2}$ at a change point), and (iii) seasonal ($d_{t}=\bar{d}+A\sin(2\pi t/T)$ for amplitude $A$).
A.2 Decision Variables
  • $x_{n,t}\geq 0$: order quantity placed by echelon $n$ in period $t$.
  • $I_{n,t}\geq 0$: on-hand inventory at echelon $n$ at the end of period $t$.
  • $B_{n,t}\geq 0$: backorders at echelon $n$ at the end of period $t$.
A.3 Objective Function
The LP minimizes total holding and backorder costs across all echelons and periods:
$\text{minimize}\quad\sum_{n\in\mathcal{N}}\sum_{t\in\mathcal{T}}\bigl(h_{n}\,I_{n,t}+b_{n}\,B_{n,t}\bigr).$ (9)
A.4 Constraints
C1: Inventory balance (retail echelon).
For the retailer ($n=1$):
$I_{1,t}-B_{1,t}=I_{1,t-1}-B_{1,t-1}+x_{1,t-L_{1}}-d_{t},\quad\forall\,t\in\mathcal{T},$ (10)
where $x_{1,t-L_{1}}$ represents the order placed $L_{1}$ periods ago that arrives in period $t$. For $t-L_{1}\leq 0$, we set $x_{1,t-L_{1}}=0$ (no arrivals from pre-horizon orders unless initial pipeline inventory is modeled separately).
C2: Inventory balance (upstream echelons).
For each non-retail echelon $n\in\{2,\ldots,N\}$:
$I_{n,t}-B_{n,t}=I_{n,t-1}-B_{n,t-1}+x_{n,t-L_{n}}-\sum_{c\in\text{children}(n)}x_{c,t},\quad\forall\,t\in\mathcal{T},$ (11)
where $\text{children}(n)=\{n{-}1\}$ in the serial topology. The demand faced by echelon $n$ is the order placed by its downstream neighbor, inducing the propagation mechanism that generates the bullwhip effect (Lee et al., 1997; Chen et al., 2000).
C3: Capacity constraints.
$x_{n,t}\leq C_{n},\quad\forall\,n\in\mathcal{N},\;t\in\mathcal{T}.$ (12)
C4: Non-negativity.
$x_{n,t},\;I_{n,t},\;B_{n,t}\geq 0,\quad\forall\,n\in\mathcal{N},\;t\in\mathcal{T}.$ (13)
C5: Initial conditions.
$I_{n,0}=I_{n,0}^{\text{init}},\quad B_{n,0}=0,\quad\forall\,n\in\mathcal{N}.$ (14)
C6: Demand satisfaction (implicit).
Backorders $B_{1,t}$ at the retail echelon capture unmet demand. The inventory balance constraint (10) ensures that $B_{1,t}\geq d_{t}-I_{1,t-1}-x_{1,t-L_{1}}+B_{1,t-1}$ whenever $I_{1,t}=0$ (i.e., stock is depleted). The backorder cost $b_{1}$ in the objective penalizes unmet demand, so the LP trades off between holding excess inventory and incurring backorder penalties.
Appendix B Error Type Specifications
This appendix defines the injection mechanism and solver signature for each of the ten error types introduced in Section 3.2. While the main text groups errors by the operational subsystem they affect, Table 7 documents how each error is injected into a valid LP and how it manifests in Gurobi’s IIS output. The difficulty column reflects empirical solve rates: “Hard” errors (ME-1, ME-2, ME-4, ME-10) generate misleading IIS signals that require multi-echelon causal reasoning; “Medium” errors produce localized but ambiguous signals; “Easy” errors (ME-3, ME-5) yield IIS patterns that directly pinpoint the violated constraint or produce feasible-but-irrational solutions detected by the oracle.
Table 7: Error type injection specifications.
Type Injection Mechanism IIS Signature Difficulty
[…]
ME-2 Remove arrival term from all inventory balance constraints at target echelon Multiple inv_balance constraints + backorder caps Hard
ME-3 Force net inventory ($I_{n,t}-B_{n,t}$) to value below achievable minimum Single inv_balance constraint with impossible target Easy
[…]
ME-6 Add constraints forcing $x_{n,t}\geq x_{n-1,t-1}+\text{offset}$ exceeding capacity Multiple bullwhip_force constraints + supply cap Medium
ME-7 Perturb arrival coefficient from 1.0 to 0.050.050.20.2 in inventory balance Multiple inv_balance constraints with modified coefficient Medium
ME-8 Flip sign in demand propagation ($d_{n,t}=-x_{n-1,t}$ instead of ++) Multiple demand_prop constraints (negated term) Medium
ME-9 Add minimum-order constraints at factory exceeding supply capacity Multiple min_order constraints + supply cap Medium
[…]
Appendix C Phase II Rationality Checks: Definitions and Thresholds
The Phase II RationalityOracle evaluates whether a feasible LP solution is operationally rational. It consists of five domain-specific checks, each derived from supply chain theory. Not all checks apply to every error type: the oracle uses an error-type-aware mapping to avoid penalizing correct repairs for artifacts introduced by the error injection process.
C.1 Five Rationality Checks
Check 1: Base-stock structure.
For each echelon $n$, the oracle verifies that inventory levels approximate the echelon base-stock policy established by Clark and Scarf (1960). The check computes the coefficient of variation of inventory:
$\text{CV}_{n}=\frac{\text{Std}(I_{n,t})}{\text{Mean}(I_{n,t})},$ (15)
and flags a violation when $\text{CV}_{n}>\tau_{\text{bs}}$ with $\tau_{\text{bs}}=2.0$. A well-managed inventory under a base-stock policy exhibits relatively stable levels (low CV), while a poorly repaired model may produce wild inventory fluctuations.
Check 2: Bullwhip ratio.
The bullwhip effect (Lee et al., 1997) measures order variance amplification. For each upstream echelon $n\geq 2$, the oracle computes:
$\text{BWR}_{n}=\frac{\text{Var}(x_{n,t})}{\text{Var}(d_{t})},$ (16)
where $d_{t}$ is the external customer demand. The check flags a violation when $\text{BWR}_{n}>\tau_{\text{bw}}$ with $\tau_{\text{bw}}=3.0$. The threshold is set above the theoretical lower bound of Chen et al. (2000) (which depends on lead time and forecasting method) to accommodate the LP’s deterministic setting.
Check 3: Inventory allocation.
In a well-structured serial system, inventory should not concentrate disproportionately at the retailer while upstream echelons hold nothing. The oracle flags a violation when the retailer’s average on-hand inventory exceeds all upstream echelons combined and is meaningful relative to demand:
$\bar{I}_{1}>\tau_{\text{ia}}\cdot\bar{d}\;\;\text{and}\;\;\sum_{n=2}^{N}\bar{I}_{n}\approx 0,$ (17)
where $\bar{I}_{n}=\frac{1}{T}\sum_{t}I_{n,t}$ and $\tau_{\text{ia}}=0.25$. This check detects structural cost misallocation—particularly cost-swap errors (ME-5) where the optimizer places inventory at the wrong echelon because holding cost coefficients are distorted.
Check 4: Cost consistency.
The oracle enforces two properties of the cost structure. First, holding costs in the problem configuration should decrease upstream: $h_{1}\geq h_{2}\geq\cdots\geq h_{N}$ (with a 1% tolerance). Second, if the Gurobi model is available, the oracle checks that each variable’s objective coefficient matches the configuration—that is, the model’s holding and backorder cost coefficients equal the values specified in the problem parameters. This check detects repairs that corrupt cost coefficients or tamper with the objective function.
Check 5: Order smoothing.
The oracle checks that order quantities do not exhibit extreme oscillation:
$\frac{\max_{t}|x_{n,t}-x_{n,t-1}|}{\text{Mean}(x_{n,t})}<\tau_{\text{os}},$ (18)
with $\tau_{\text{os}}=5.0$. Extreme order oscillation (the “nervous system” problem in MRP systems) indicates an unstable replenishment policy that would be operationally infeasible in practice due to production changeover costs and capacity constraints.
C.2 Which Checks Apply to Which Error Types
Table 8 specifies which checks apply to each error type. Excluded checks are those where the saboteur’s error injection or the associated constraint tightening creates structural conditions that cause the check to fail even on correct repairs.
Table 8: Oracle check applicability by error type. A check mark (✓) indicates the check is applied; a dash (–) indicates exclusion. The “Exclusion rationale” column explains why excluded checks produce false failures for that error type.
Error
Base-stock
Bullwhip
Inv. alloc.
Cost consist.
Order smooth.
Exclusion rationale
ME-1 Demand inflation creates legitimate order variance
ME-2 Lead-time repair creates legitimate order variance
ME-3 Balance repair creates legitimate order variance
ME-4 Capacity change creates legitimate order variance
ME-5 Cost-only error; cost consistency detects the root cause
ME-6 Bullwhip error; only order smoothing is artifact-free
ME-7 Coefficient error; cost consistency detects tampering
ME-8 Sign error; cost consistency detects coefficient changes
ME-9 Constraint removal creates legitimate order variance
ME-10 Index correction creates legitimate order variance
Appendix D Prompt and Interaction Examples
This appendix presents the full system prompt used by OptiRepair agents and a complete interaction transcript illustrating both phases of the closed-loop pipeline.
D.1 System Prompt
The following system prompt is used for trained models (Qwen3-8B and Llama-3.1-8B). API models receive a JSON-formatted variant with identical content.
System Prompt You are an expert supply chain optimization debugger. Your task is to debug infeasible multi-echelon supply chain LP models through a two-phase closed-loop process. == Two-Phase Process == Phase 1 (Debug): Fix infeasibility. Identify conflicting constraints and repair them to restore feasibility. Phase 2 (Validate): After restoring feasibility, an automated rationality oracle checks the solution for operational rationality. A solver-feasible solution is not necessarily a good supply chain policy. Checks: 1. Base-stock rationality -- cost-efficient inventory
2. Bullwhip control -- order variance within bounds
3. Monotonicity -- no severe inventory inversions
4. Cost consistency -- coefficients match specification
5. Order smoothing -- no excessive order oscillation
If the solution fails, you receive violation feedback and must fix the structural issue. == Available Actions == - GET_IIS(): Compute Irreducible Infeasible Subsystem
- DROP_CONSTRAINT(name): Remove a constraint
- RELAX_CONSTRAINT(name, amount): Relax a constraint
- UPDATE_RHS(name, value): Update right-hand side
- UPDATE_OBJ(var_prefix, value): Update objective coeff
- UPDATE_BOUNDS(var, lb, ub): Update variable bounds
- CHECK_SLACK(name): Check slack for a constraint
- SUBMIT(): Submit the repaired model
== Response Format == <think>
Your reasoning about the infeasibility or
rationality violation...
</think>
Action: ACTION_TYPE(arguments)
D.2 State Observation Format
At each step, the agent receives a structured observation containing the solver status, IIS information (if applicable), model dimensions, and recent action history.
State Observation ## Current State
- Solver Status: INFEASIBLE
- Step: 0
## IIS (Irreducible Infeasible Subsystem)
- Conflicting Constraints: [’capacity_e1_t1’,
 ’capacity_e1_t3’, ’capacity_e1_t5’,
 ’demand_balance_e1_t1’, ’demand_balance_e1_t3’]
- Conflicting Bounds: [’x_e1_t1 <= 20’,
 ’x_e1_t3 <= 20’, ’x_e1_t5 <= 20’]
## Model Structure
- Total Constraints: 142
- Total Variables: 96
- Constraint Names (first 10): [’demand_balance_e1_t1’,
 ’demand_balance_e1_t2’, ..., ’capacity_e1_t1’, ...]
What action should be taken next?
D.3 Phase I Interaction Transcript
The following transcript shows the trained Qwen3-8B model debugging an ME-4 error (capacity too tight) on a 3-echelon, 16-period supply chain.
Agent – Turn 1 <think>
The model is infeasible. The IIS shows capacity constraints at echelon 1 (capacity_e1_t1, t3, t5) conflicting with demand balance constraints. The variable bounds confirm x_e1 <= 20 for all flagged periods. Average demand is 45 units, well above the capacity of 20. This is an ME-4 error: capacity set far below demand, causing cascading infeasibility. I should relax the echelon-1 capacity constraints.
</think>
Action: RELAX_CONSTRAINT(capacity_e1, 50.0)
Environment Response ## Current State
- Solver Status: OPTIMAL
- Step: 1
- Objective Value: 12487.3
-> Entering Phase 2 (Rationality Oracle)
D.4 Phase II Interaction Transcript
After Phase I restores feasibility, the rationality oracle evaluates the solution.
Rationality Oracle – Feedback ## Rationality Feedback (from Phase 2 Oracle) Echelon 1 (Retailer, h=\$8.00) holds 45.2 units on-hand inventory while Echelon 2 (Warehouse, h=\$3.00) holds 12.3 units. Holding cost at Retailer (\$8.00) > Warehouse (\$3.00). Cost-efficient allocation violated. The solver found a feasible solution, but it failed rationality checks. The repair restored feasibility but created an operationally irrational policy. Fix the structural issue described above, then restore feasibility again. ## Closed-Loop Status
- Phase: VALIDATE -> DEBUG (loop-back)
- Debug-Validate Loop: 1
Agent – Turn 2 <think>
The capacity relaxation fixed infeasibility but broke cost consistency. The retailer (echelon 1) has holding cost h=\$8 but holds 45 units, while the warehouse (echelon 2) with h=\$3 holds only 12 units. In a rational supply chain, the cheaper storage location should buffer more inventory. The holding cost coefficient for echelon 1 in the objective function may be distorted. I should update it to match the warehouse’s cost structure.
</think>
Action: UPDATE_OBJ(hold_e1, 3.0)
Environment Response ## Current State
- Solver Status: OPTIMAL
- Step: 2
- Objective Value: 11832.1
-> Re-checking Rationality Oracle
Rationality Oracle – Result ## Rationality Check: PASSED
- Base-stock rationality: PASS
- Bullwhip control: PASS
- Monotonicity: PASS
- Cost consistency: PASS
- Order smoothing: PASS
Episode complete. Solution is both solver-feasible and operationally rational.
Total steps: 2 | Loops: 1
D.5 API Model Response Format
API models (GPT-5.2, Claude, Gemini, etc.) use a JSON response format instead of the <think>/Action: format:
API Agent Response (JSON) {
 "reasoning": "The IIS shows capacity constraints
    at echelon 1 conflicting with demand balance.
    Capacity C1=20 is below demand of 45. I should
    relax the capacity constraint.",
 "action": "RELAX_CONSTRAINT",
 "target": "capacity_e1",
 "value": 50.0
Both formats are parsed into identical Action objects by the environment’s action parser. The trained model format is more token-efficient (no JSON overhead) and allows natural chain-of-thought reasoning within the <think> tags.
Appendix E Full Per-Error-Type Results
This appendix reports the complete per-error-type breakdown for all models evaluated on OptiRepair-SC. Figure 7 in the main text decomposes overall performance into Phase 1 Recovery Rate and Phase 2 Pass Rate; the tables below give the per-error-type detail. Table 9 presents per-error-type RRR for both trained models and the top-6 API baselines. Table 10 presents the complete RRR breakdown for all 22 API models in landscape format.
Table 9: Per-error-type RRR (%) on OptiRepair-SC for trained models and top-6 API baselines. ME-5 (cost structure) is universally solvable; ME-2 (lead time) and ME-4 (capacity) separate trained models from API baselines.
Model ME-1 ME-2 ME-3 ME-4 ME-5 ME-6 ME-7 ME-8 ME-9 ME-10
Qwen3-8B (trained) 85.2 64.5 96.7 61.3 100 89.3 91.7 88.0 96.4 86.7
Llama-3.1-8B (trained) 81.5 67.7 93.3 74.2 100 85.7 87.5 84.0 92.9 83.3
GPT-5.2 7.4 41.9 73.3 38.7 100 7.1 50.0 36.0 57.1 6.7
Gemini 2.5 Pro 0.0 6.5 63.3 6.5 100 60.7 50.0 60.0 57.1 23.3
DeepSeek-R1 3.7 12.9 76.7 9.7 100 0.0 58.3 60.0 42.9 10.0
Gemini 2.5 Flash 0.0 3.2 66.7 9.7 100 21.4 41.7 44.0 21.4 13.3
Claude Opus 4.5 3.7 6.5 70.0 3.2 100 35.7 37.5 20.0 25.0 0.0
Qwen2.5-7B 40.7 0.0 66.7 9.7 90.0 3.6 29.2 36.0 0.0 33.3
Table 10: RRR (%) by error type for all 22 API models and the trained Qwen3-8B pipeline on OptiRepair-SC test set. Bold = highest API score per error type.
Rank Model ME-1 ME-2 ME-3 ME-4 ME-5 ME-6 ME-7 ME-8 ME-9 ME-10 Overall
1 GPT-5.2 7.4 41.9 73.3 38.7 100 7.1 50.0 36.0 57.1 6.7 42.2
2 Gemini 2.5 Pro 0.0 6.5 63.3 6.5 100 60.7 50.0 60.0 57.1 23.3 42.2
3 DeepSeek-R1-0528 3.7 12.9 76.7 9.7 100 0.0 58.3 60.0 42.9 10.0 37.0
4 Gemini 2.5 Flash 0.0 3.2 66.7 9.7 100 21.4 41.7 44.0 21.4 13.3 32.0
5 Qwen2.5-7B 40.7 0.0 66.7 9.7 90.0 3.6 29.2 36.0 0.0 33.3 31.0
6 Claude Opus 4.5 3.7 6.5 70.0 3.2 100 35.7 37.5 20.0 25.0 0.0 30.3
7 Qwen2.5-Max 18.5 0.0 63.3 9.7 100 0.0 50.0 16.0 0.0 10.0 26.8
8 GPT-o4-mini 7.4 0.0 50.0 0.0 100 7.1 25.0 16.0 39.3 10.0 25.7
9 Claude Sonnet 4 3.7 6.5 66.7 3.2 100 0.0 12.5 8.0 25.0 3.3 23.6
10 DeepSeek-V3.2 3.7 0.0 50.0 3.2 100 10.7 12.5 20.0 14.3 0.0 21.8
11 Gemini 2.0 Flash 0.0 0.0 80.0 0.0 100 0.0 8.3 0.0 3.6 0.0 20.1
12 Claude Haiku 4.5 0.0 0.0 23.3 0.0 100 7.1 25.0 28.0 7.1 0.0 19.0
13 GPT-o3 0.0 0.0 36.7 3.2 100 0.0 8.3 0.0 32.1 3.3 19.0
14 GPT-4.1 0.0 3.2 36.7 6.5 100 0.0 8.3 4.0 10.7 0.0 17.6
15 GPT-4.1 Mini 3.7 0.0 26.7 0.0 100 0.0 4.2 0.0 32.1 0.0 17.2
16 GPT-5 Mini 0.0 0.0 33.3 0.0 100 0.0 12.5 0.0 14.3 3.3 16.9
17 GPT-o1 0.0 0.0 20.0 0.0 96.7 0.0 4.2 0.0 7.1 0.0 13.5
18 Claude 3.7 Sonnet 0.0 0.0 16.7 0.0 100 0.0 8.3 0.0 3.6 0.0 13.4
19 Qwen2.5-14B 0.0 0.0 16.7 0.0 100 0.0 4.2 0.0 0.0 0.0 12.7
20 Qwen2.5-32B 0.0 0.0 16.7 0.0 20.0 0.0 0.0 0.0 0.0 3.3 4.2
21 Kimi K2 0.0 0.0 3.3 0.0 10.0 0.0 0.0 0.0 0.0 0.0 1.4
22 Llama-3.3-70B 0.0 0.0 0.0 0.0 3.3 0.0 0.0 0.0 0.0 0.0 0.4
Trained Qwen3-8B 85.2 64.5 96.7 61.3 100 89.3 91.7 88.0 96.4 86.7 81.7
Appendix F Guided Prompt Experiments
We evaluated two guided prompt variants on Claude Opus 4.5 and GPT-5.2-chat. The anti-loop variant added explicit instructions to avoid repeating failed repair strategies; the domain-knowledge variant injected supply chain structural hints (base-stock properties, bullwhip ratio interpretation, cost structure constraints). The domain-knowledge prompt improved Opus by 7.4 percentage points (RRR 30.3% to 37.7%) but degraded GPT-5.2 by 24.1 percentage points (42.2% to 18.1%). The anti-loop prompt hurt both models. Prompt text is available from the authors upon request.
Appendix G Dataset Construction Details
This appendix describes the generation pipeline, instance parameters, validation procedures, and dataset statistics for OptiRepair-SC.
G.1 OptiRepair-SC Generation Pipeline
We construct the OptiRepair-SC problem set through a four-stage pipeline:
  1. 1.
    Source instance generation. We generate supply chain LP instances by sampling parameters from the ranges specified in Table 11. Each instance is solved to optimality by Gurobi 12.0 to confirm a feasible baseline. Trivially feasible instances (optimal objective value below a threshold) or degenerate ones (fewer than 10 active constraints at optimality) are filtered out.
  2. 2.
    Sabotage. For each valid source instance, the saboteur selects an error type (ME-1ME-10) and applies the corresponding injection mechanism (Appendix B). The saboteur first tightens the model with operational bounds—backorder caps and factory supply caps calibrated to the optimal solution—to prevent the LP’s natural slack from absorbing the perturbation. The error is then injected into the tightened model.
  3. 3.
    Gurobi verification. Each sabotaged model is solved by Gurobi. The model must return Infeasible status (or Optimal-but-irrational for ME-5), and the IIS computation must return a non-empty subsystem. Models that fail either check are discarded.
  4. 4.
    Oracle labeling. The ground-truth fix is applied, and the repaired model is re-solved. The solution must achieve Optimal status and pass all applicable RationalityOracle checks (Appendix C). Problems where the ground-truth fix does not pass all checks are discarded.
G.2 Instance Parameters
Table 11: Parameter ranges for OptiRepair-SC instance generation.
Parameter Range Distribution
Number of echelons NN $\{2,3,4,5\}$ Uniform
Planning periods $T$ $\{12,16,20,24\}$ Uniform
Topology Serial Fixed
Holding cost $h_{n}$ $[1,10]$ per unit/period Uniform
Backorder cost $b_{n}$ $[5,50]$ per unit/period Uniform
$h$-to-$b$ ratio 1:21{:}2 to 1:101{:}10 Derived
Capacity $C_{n}$ $[50,500]$ units/period Uniform
Lead time $L_{n}$ $\{1,2,3\}$ periods Uniform
Initial inventory $I_{n,0}^{\text{init}}$ $0,2\bar{d}]$ Uniform
Mean demand $\bar{d}$ $[50,200]$ units/period Uniform
Demand pattern Stationary, step-change, seasonal Categorical (equal)
G.3 Dataset Statistics
Table 12: Dataset split statistics for OptiRepair-SC.
Train Test Total
Total problems 692 284 976
ME-1 (Demand Inflation) 78 27 105
ME-2 (Lead Time Error) 89 31 120
ME-3 (Balance Violation) 87 30 117
ME-4 (Capacity Reduction) 89 31 120
ME-5 (Cost Structure Error) 71 30 101
ME-6 (Bullwhip Amplification) 40 28 68
ME-7 (Coefficient Perturbation) 66 24 90
ME-8 (Sign Error) 71 25 96
ME-9 (Redundant Constraint) 56 28 84
ME-10 (Index Mismatch) 45 30 75
Avg. variables per instance 185
Avg. constraints per instance 162
Avg. IIS size 12.3 constraints
The train/test split is stratified by error type, with disjoint supply chain instances (i.e., no source instance appears in both splits). The test set contains 284 problems, providing sufficient statistical power to detect differences of 5 percentage points or larger at 95% confidence.